OpenStudy (anonymous):

In my math book, it tells me to use some method with +'s and -'s to solve inequalities. Can somebody explain this to me?

OpenStudy (anonymous):

help me and that's ur 500th medal.

OpenStudy (anonymous):

Solve for x: $$\frac{1}{1-x} \leq 0$$

OpenStudy (anonymous):

oops I meant $\frac{1}{1-x}\geq 0$

OpenStudy (anonymous):

$$\frac{1}{1-x}\geq 0$$

OpenStudy (anonymous):

the book teaches me like so: +/+=yes -/+=no -/-=yes what does it mean?

OpenStudy (anonymous):

That looks like the NXOR operation

OpenStudy (anonymous):

what is the NXOR operation? Can you explain it to me?

OpenStudy (anonymous):

NOT exclusive OR

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

when does the XOR operation evaluate to true?

OpenStudy (anonymous):

and what does it have to do with inequalities :-P

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

Do you know how to solve an inequality?

OpenStudy (anonymous):

sort of, but the book's way of doing it confuses me because every time I solve it the book's way, I get the exact opposite of the answer in the book

OpenStudy (anonymous):

Let's do one now:\[\text{Find all numbers x for which:}\]\[4-x < 3 - 2x\]

OpenStudy (anonymous):

How would you solve it?

OpenStudy (anonymous):

\[4-x<3-2x\] \[4-3<x\] \[1<x\]

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

oh, it's a negative x -1<x

OpenStudy (anonymous):

actually it is \[x < 1\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

Another one:\[5 - x^2 < 8\]

OpenStudy (anonymous):

5-x^2<8 5-x^2-8<0 -3-x^2<0 I got this far

OpenStudy (anonymous):

i think i need to complete the square

OpenStudy (anonymous):

There's no need to complete the square. Just move X to one side of the inequality :-D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5-x^2<8 5-x^2-8<0 -3-x^2<0 -x^2<3 x^2>-3 |x|>-3sqrt

OpenStudy (anonymous):

\[|x|>\sqrt{-3}\]

OpenStudy (anonymous):

i'm stuck

OpenStudy (anonymous):

Instead of going to \[|x|>\sqrt{-3}\] try to conclude something from \[x^2 > -3\]

OpenStudy (anonymous):

One more \[x^2-4x+3 > 0\]

OpenStudy (anonymous):

do I split it into two different equations?

OpenStudy (anonymous):

You can actually stop once you've reached the step \[x^2 > -3\]since \[x^2 \geq 0\] is true for all x.

OpenStudy (anonymous):

(x-3)(x-1)>0 the critical points are 3 and 1 x>3 and x<0

OpenStudy (anonymous):

in this situation, my book tells me to do some sort of +/- and +/+ and -/- thing

OpenStudy (anonymous):

When is the product of two numbers positive?

OpenStudy (anonymous):

when x>3

OpenStudy (anonymous):

When you multiply two positive numbers, do you get a positive answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or two negatives

OpenStudy (anonymous):

but not if one number is positive and the other is negative.

OpenStudy (anonymous):

yup, and that is not more than 0

OpenStudy (anonymous):

so in our case, we should split the problem into two situations, each with two equations

OpenStudy (anonymous):

When both numbers are positive:\[x -3 >0\ \ \ \ and\ \ \ \ x-1 >0\] When both are negative\[x-3 < 0 \ \ \ \ and\ \ \ \ x-1 < 0\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what point are you trying to make?

OpenStudy (anonymous):

There are two different solutions to the inequality

OpenStudy (anonymous):

each accounting the fact that both expressions are either both positive or negative.

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

bbut the answer is positive, right?

OpenStudy (anonymous):

not necessarily; the answer is two different inequalities

OpenStudy (anonymous):

or a compound inequality :-D

OpenStudy (anonymous):

x < 1 or x > 3

OpenStudy (anonymous):

ok,well thanks!

OpenStudy (anonymous):

that really helped me